A rational function is given. f(x)= 6x2+1/2x2+x-1

Find all vertical asymptotes, if any, of the graph of f.

Find all horizontal asymptotes, if any, of the graph of f.
Answer #1

Given 2 fal= 62²+1 2x²+x-1 To find verhcal asymptote, fut 2x²+x-1=0 22 +2a-1-150 2 x (x+1)-1(x+1)=0 (24+1) (24-11-0 25-1 or 2